# Towards Modality Generalization: A Benchmark and Prospective Analysis

### v2

@@ 4.1 Experimental Setups
Model selection methods. Differences in results due to inconsistent tuning practices can be mistakenly attributed to the algorithms under evaluation, complicating fair comparisons and masking true performance differences. This issue becomes even more pronounced in modality generalization, where the lack of a validation set that is identically distributed to the testing modality data makes model selection—such as choosing hyperparameters, training checkpoints, or architecture variants—particularly challenging. Without proper tuning practices and robust validation mechanisms, fair assessments and progress in modality generalization remain difficult to achieve. To benchmark modality generalization comprehensively and fairly, we choose a series of model selection methods, including a training-modality validation set, leave-one-modality-out cross-validation, and testing-modality validation set. We provide the details of the three model selection methods in Appendix C.
Implementation details. We describe the implementation of modal encoding, testing perceptor implementation for Strong MG, and the architecture of the learner $h$. More details like hyperparameter settings are briefly introduced in this section while being detailed in Appendix A and C.
Modal encoding. We encode the text modality using unsupervised tokenization, while treating other modalities, such as audio and depth, as visual inputs, following recent modality-binding models [30, 31, 32]. For text, we typically use a subword segmentation algorithm, such as SentencePiece [70, 71], to convert the text into token indices. RGB images are resized and cropped to 224$\times$224 pixels. Following [72] and [73], we encode 2-second audio sampled at 16kHz into spectrograms, which are 2D signals that can be processed as images. Depth data is converted into disparity maps for scale invariance. For video, we sample two frames per second and apply spatial cropping to generate sequences of 224$\times$224 images. Each image is then divided into 16 patches, as per ViT model [74].
Testing perceptors in strong modality generalization. We carefully select perceptors for the problem setting of strong modality generalization, prioritizing pre-trained models that are widely recognized within the community to facilitate training. Specifically, we use T5 [75] for the text modality [^2] https://huggingface.co/google-t5/t5-small, and ViT [74] for the other modalities [^3] https://huggingface.co/google/vit-base-patch16-224, as mentioned earlier, i.e., treating them as images. Notably, we incorporate a linear module to ensure a consistent embedding space, with 1024 as the default dimension. To adapt these models for our strong modality generalization evaluation, we retain the pre-trained weights and apply LoRA [76] with contrastive objectives. Note that we use these self-supervised methods, aiming to prevent any task-specific information leakage. We randomly mask the samples to generate two views, which are then contrasted with other samples within the same batch. Here we set the batch size to 128 for the text modality and 64 for the other modalities.
Architecture of the learner $h$. Upon pre-trained modal perceptors (i.e., modality-binding models or solely trained perceptors for testing modalities), we construct a simple learner $h$ for follow-up tasks. Specifically, we adopt a 4-layer MLP to learn task-relevant features with ReLU as an activation function, and a linear model as a classifier to transform features to class distributions. The details are demonstrated in Table 3, Appendix A.
We adapt various algorithms while maintaining consistency in their model architecture and optimization strategies. The model selection methods are implemented following DomainBed and detailed in Appendix C. To ensure a thorough comparison, we perform an extensive hyperparameter search, randomly selecting hyperparameters and random seeds. For each algorithm, we conduct 9 trials over hyperparameter distributions (see Table 4) with 3 different random seeds, considering parameters such as learning rate, weight decay, batch size, and other specialized hyperparameters. Refer to more details in Appendix A.
@@
Figure 3: Weak MG performance comparison for diverse algorithms with three perceptors (“IB”$\leftrightarrow$ ImageBind, “LB”$\leftrightarrow$ LanguageBind, and “UB”$\leftrightarrow$ UniBind). The dashed line indicates the averaged performance for two categories of algorithms (MML and DG) w.r.t. different perceptors. The bar represents the average performance of each algorithm across different modalities.
@@
Figure 4: Weak MG performance comparison for different model selection methods across different benchmarks.
@@
Figure 5: Weak MG Performance comparison for specific testing modalities across different algorithms.
4.2 Experimental Results of Weak Modality Generalization
@@
The empirical results under the setting of weak modality generalization are shown in Figures 35, and 5. Here we report the testing results for each dataset across three modal selection methods. The results on different modalities are detailed in Tables 5, 6, and 7, Appendix D. From these results, we derive the following findings.
@@
Performance gap between MML and DG algorithms. As demonstrated in Figure 3, the comparison between MML and DG shows that DG outperforms MML for weak modality generalization. MML typically focuses on joint embedding across modalities. This makes MML effective at leveraging the complementary information within the modalities, potentially achieving better performance in in-modal tasks. However, in the case of weak modality generalization where the model encounters unseen modalities during testing, MML struggles to generalize effectively. In contrast, DG is specifically designed to improve generalization across different domains, and its performance remains more robust when tested on unseen modalities. As a result, DG is better suited for weak modality generalization compared to MML.
@@
Figure 6: Strong MG performance comparison for diverse algorithms with three perceptors. The dashed line indicates the averaged performance for different perceptors w.r.t. two categories of algorithms (MML and DG). The gray line represents the performance of a random prediction. The bar indicates the average performance across different testing modalities.
@@
Figure 7: Strong MG performance comparison for different model selection methods across different benchmarks.
@@
Figure 8: Strong MG Performance comparison for specific testing modalities across different algorithms.
@@
Better modality binding $\neq$ better generalization. Better modality binding does not necessarily translate to improved performance in modality generalization. For example, LanguageBind, despite achieving superior performance in modality binding, underperforms in most cases compared to the other two modality-binding methods under the Weak MG setting. This highlights the inherent limitations of learning a joint embedding space for modality generalization, as discussed in Section 2.2. Furthermore, this distinction manifests the core difference in learning objectives between modality binding and generalization. Modality binding focuses on identifying the shared information across existing modalities or around specific modalities (e.g., image or language). In contrast, modality generalization seeks to capture invariants across all modalities, including novel ones. These differing objectives lead to inconsistencies. This finding calls for further exploration and development of methods specifically tailored for modality generalization.
@@
The importance of model selection. The performance varies significantly across different model selection methods, as shown in Figure 5. This is consistent with findings from DomainBed [52], which demonstrate that validation assumptions influence model selection criteria and ultimately determine modality generalization performance during testing. Among the three model selection methods, the testing-modality validation set (oracle) approach outperforms the others overall. By selecting the model that achieves the best performance on the validation set derived from the test modality, this method thereby optimizes the model to generalize across unseen modalities. These results emphasize that precise model selection is critical to the success of the experiment.
@@
Modality differences for generalization. We demonstrate the performance across different testing modalities, as depicted in Figure 5. A clear prioritization of modalities is observed across different algorithms. For instance, generalization to “Lan” is notably easier on both NYUDv2 and VGGSound-S. In the Weak MG setting, existing perceptors are incorporated for their joint embedding spaces, which are typically designed for semantic purposes (e.g., CLIP [13] leverages natural language to reference learned visual concepts, enhancing the development of recent modality-binding methods [30, 31, 32]).
4.3 Experimental Results of Strong Modality Generalization
@@
We present the evaluation results under the strong modality generalization setting in Figures 68, and 8 (details in Tables 89, and 10, Appendix D). Unlike the weak modality generalization setting, the testing perceptors are exclusively trained on data from novel modalities by self-supervised learning. From these results, we derive the following observations.
@@
Narrowed performance gap between MML and DG. In the strong modality generalization setting, the performance gap between MML and DG methods diminishes, with MML slightly outperforming DG, especially on the NYUDv2 dataset. This contrasts with the previous weak setting, where DG methods hold a significant advantage over MML, as shown in Figure 6. Notably, MML demonstrates more robust performance across various hyperparameter searches and model selection strategies. This can be attributed to its learning objective that focuses on exploiting the complementarity among multi-modalities, not aligning well with the modality generalization goal. As a comparison, DG focuses on domain-invariant features (i.e., modal-invariant in our case), which are more sensitive when testing modalities are completely isolated.
@@
Model selection remains crucial. The performance in the strong modality generalization setting still varies with different model selection methods, highlighting the importance of model selection in Figure 8. Without a validation set that mirrors the test modality, hyperparameter tuning and performance benchmarking become particularly challenging. Among the three methods, the training-modality validation set leads to more stable performance across algorithms compared to the other two model selections. The phenomenon is different from that in the case of weak modality generalization. This inspires researchers to design task-specific model selection methods or more general algorithms that can consistently perform well across multiple modality generalization approaches.
@@
Limited utility of modality binding. Modality binding methods improve the generalization in weak modality generalization, while exhibiting limited utility in a more challenging task, strong modality generalization. The alignment between the training and testing perceptors helps the model preserve the relationship between modalities, leading to better consistency and accuracy when processing novel modalities in weak modality generalization testing. However, their utility diminishes in strong modality generalization, where the testing modality’s perceptor is isolated and does not share alignment with training modalities. This is due to their reliance on a joint embedding space, which struggles to generalize to unseen modalities with isolated perceptors. These findings suggest a need for alternative strategies that move beyond binding-centric paradigms to effectively handle Strong MG scenarios.
@@
Narrowed generalization gap between modalities. In the Strong MG setting, the difficulty of generalizing to the testing modality becomes more monotonic, as shown in Figure 8. With the limited utility of modality binding, the distinction of “easily generalizable” modalities is reduced. However, performance remains higher when “Lan” is the testing modality, particularly on NYUDv2 and VGGSound-S. Despite complete isolation of testing perceptors, connections between modalities can still be identified, which enhances generalization and substantiates our motivation to embrace Strong MG.
@@
Potential of strong modality generalization. Although the overall performance in strong modality generalization is lower than in weak modality generalization (see Figure 6), all methods achieve better-than-random predictions, indicating the ability to capture some cross-modal invariants. However, the performance gap between weak and strong modality generalization emphasizes the complexity of the strong modality generalization task, underlining the need for models that generalize without tightly bound perceptors or shared embedding spaces, such as those enhanced by self-supervised perceptors for modality generalization.
5 Discussions and Outlook
@@
Modality generalization beyond domain generalization. Modality generalization (MG) extends beyond the traditional scope of domain generalization by focusing on enabling models to generalize across entirely unseen modalities rather than unseen domains. While domain generalization assumes similar underlying features across domains, MG introduces additional complexity due to the vast heterogeneity between modalities, such as text, audio, and visual data. Future research should emphasize developing strategies to learn and transfer invariant features across such diverse modalities. This includes moving beyond shared embedding spaces to identify deeper, task-agnostic representations that are transferable across unseen modalities. To address these challenges, researchers could explore the following aspects: (1) cross-modal invariant learning: develop mechanisms to extract invariants that are robust across modalities without assuming tight alignment or correspondence; (2) hybrid generalization frameworks: combine domain-specific techniques (e.g., meta-learning) with modality-aware methods to enhance adaptability and robustness; (3) realistic scenario validation: test models under resource-constrained or privacy-sensitive scenarios where data availability is limited for certain modalities.
@@
Specific algorithm design for modality generalization. Designing algorithms specifically tailored to modality generalization is critical for addressing the limitations of existing multi-modal learning frameworks. First, leveraging and advancing self-supervised learning to create perceptors for unseen modalities can enable effective feature extraction even in the absence of prior alignments. Second, exploring dynamic modality binding mechanisms that adapt to the availability of training modalities and the characteristics of unseen testing modalities can enhance model flexibility. Finally, developing task-aware modality generalization algorithms that integrate task-specific priors can improve generalization for downstream applications.
@@
Construction of related datasets. The lack of comprehensive datasets for MG research remains a major challenge. Future datasets should include a diverse range of modalities, such as text, audio, video, depth, and bio-signals, to better simulate the complexity of real-world scenarios. These datasets should be designed to reflect MG-specific challenges, intentionally excluding certain modalities during training while ensuring their presence in the testing set to create both weak and strong MG scenarios. To facilitate benchmarking for large-scale models, datasets should also vary in complexity, modality richness, and data volume, addressing scalability concerns. Additionally, real-world constraints such as privacy and resource limitations should be considered, incorporating sensitive modalities like medical imaging or encrypted communications to emphasize the practicality of MG use cases.

### v3

@@ 5.4. Model Selection Methods
Differences in results due to inconsistent tuning practices can be mistakenly attributed to the algorithms under evaluation, complicating fair comparisons and masking true performance differences. This issue becomes even more pronounced in modality generalization, where the lack of a validation set that is identically distributed to the testing modality data makes model selection, such as choosing hyperparameters, training checkpoints, or architecture variants, particularly challenging. Without proper tuning practices and robust validation mechanisms, fair assessments and progress in modality generalization remain difficult to achieve. To benchmark modality generalization comprehensively and fairly, we choose a series of model selection methods, including training-modality validation (TM), leave-one-modality-out cross-validation (LOO), and testing-modality validation (Oracle).
@@
Figure 3. Weak MG performance comparison for diverse algorithms with three perceptors (“IB”$\leftrightarrow$ ImageBind, “LB”$\leftrightarrow$ LanguageBind, and “UB”$\leftrightarrow$ UniBind). The dashed line indicates the averaged performance for two categories of algorithms (MML and DG) w.r.t. different perceptors. The bar represents the average performance of each algorithm across different modalities.
@@
Figure 4. Averaged performance comparison for different model selection methods across different benchmarks (Weak MG).
@@
Figure 5. Weak MG Performance comparison for specific testing modalities across different algorithms.
5.5. Implementation Details
We describe the implementation of modal encoding, testing perceptor implementation for Strong MG, and the architecture of the learner $h$.
$\bullet$ Modal encoding. We encode the text modality using unsupervised tokenization, while treating other modalities, such as audio and depth, as visual inputs, following recent modality-binding models (Girdhar et al., 2023; Zhu et al., 2024; Lyu et al., 2024). For text, we typically use a subword segmentation algorithm, such as SentencePiece (Sennrich, 2016; Kudo, 2018), to convert the text into token indices. RGB images are resized and cropped to 224$\times$224 pixels. Following (Gong et al., 2021) and (Girdhar et al., 2022), we encode 2-second audio sampled at 16kHz into spectrograms, which are 2D signals that can be processed as images. Depth data is converted into disparity maps for scale invariance. For video, we sample two frames per second and apply spatial cropping to generate sequences of 224$\times$224 images. Each image is then divided into 16 patches, as per ViT model (Dosovitskiy et al., 2021).
$\bullet$ Testing perceptors in strong modality generalization. We carefully select perceptors for the problem setting of strong modality generalization, prioritizing pre-trained models that are widely recognized within the community to facilitate training. Specifically, we use T5 (Raffel et al., 2020) for the text modality [^2] https://huggingface.co/google-t5/t5-small, and ViT (Dosovitskiy et al., 2021) for the other modalities [^3] https://huggingface.co/google/vit-base-patch16-224, as mentioned earlier, i.e., treating them as images. Notably, we incorporate a linear module to ensure a consistent embedding space, with 1024 as the default dimension. To adapt these models for our strong modality generalization evaluation, we retain the pre-trained weights and apply LoRA (Hu et al., 2021) with contrastive objectives. Note that we use these self-supervised methods, aiming to prevent any task-specific information leakage. We randomly mask the samples to generate two views, which are then contrasted with other samples within the same batch. Here we set the batch size to 128 for the text modality and 64 for the other modalities.
$\bullet$ Architecture of the learner $h$. Upon pre-trained modal perceptors (i.e., modality-binding models or solely trained perceptors for testing modalities), we construct a simple learner $h$ for follow-up tasks. Specifically, we adopt a 4-layer MLP to learn task-relevant features with ReLU as an activation function, and a linear model as a classifier to transform features to class distributions.
We adapt various algorithms while maintaining consistency in their model architecture and optimization strategies. The model selection methods are implemented following DomainBed. To ensure a thorough comparison, we perform an extensive hyperparameter search, randomly selecting hyperparameters and random seeds. For each algorithm, we conduct 9 trials over hyperparameter distributions (see Table 4) with 3 different random seeds, considering parameters such as learning rate, weight decay, batch size, and other specialized hyperparameters.
6. Results & Analysis
We elaborate on the results and corresponding analysis for two types of modality generalization in this section.
6.1. Results of Weak Modality Generalization
@@
The empirical results under the setting of weak modality generalization are shown in Figures 35, and 5. Here we report the testing results for each dataset across three modal selection methods. From these results, we derive the following findings.
@@
$\bullet$ Performance gap between MML and DG algorithms. As demonstrated in Figure 3, the comparison between MML and DG shows that DG outperforms MML for weak modality generalization. MML typically focuses on joint embedding across modalities. This makes MML effective at leveraging the complementary information within the modalities, potentially achieving better performance in in-modal tasks. However, in the case of weak modality generalization where the model encounters unseen modalities during testing, MML struggles to generalize effectively. In contrast, DG is specifically designed to improve generalization across different domains, and its performance remains more robust when tested on unseen modalities. As a result, DG is better suited for weak modality generalization compared to MML.
@@
Figure 6. Strong MG performance comparison for diverse algorithms with three perceptors. The dashed line indicates the averaged performance for different perceptors w.r.t. two categories of algorithms (MML and DG). The gray line represents the performance of a random prediction. The bar indicates the average performance across different testing modalities.
@@
Figure 7. Averaged performance comparison for different model selection methods across different benchmarks (Strong MG).
@@
Figure 8. Strong MG Performance comparison for specific testing modalities across different algorithms.
@@
$\bullet$ Better modality binding $\neq$ better generalization. Better modality binding does not necessarily translate to improved performance in modality generalization. For example, LanguageBind, despite achieving superior performance in modality binding, underperforms in most cases compared to the other two modality-binding methods under the Weak MG setting. This highlights the inherent limitations of learning a joint embedding space for modality generalization, as discussed in Section 3.4. Furthermore, this distinction manifests the core difference in learning objectives between modality binding and generalization. Modality binding focuses on identifying the shared information across existing modalities or around specific modalities (e.g., image or language). In contrast, modality generalization seeks to capture invariants across all modalities, including novel ones. These differing objectives lead to inconsistencies. This finding calls for further exploration and development of methods specifically tailored for modality generalization.
@@
$\bullet$ The importance of model selection. The performance varies significantly across different model selection methods, as shown in Figure 5. This is consistent with findings from DomainBed (Gulrajani and Lopez-Paz, 2021), which demonstrate that validation assumptions influence model selection criteria and ultimately determine modality generalization performance during testing. Among the three model selection methods, the testing-modality validation set (oracle) approach outperforms the others overall. By selecting the model that achieves the best performance on the validation set derived from the test modality, this method thereby optimizes the model to generalize across unseen modalities. These results emphasize that precise model selection is critical to the success of the experiment.
@@
$\bullet$ Modality differences for generalization. We demonstrate the performance across different testing modalities, as depicted in Figure 5. A clear prioritization of modalities is observed across different algorithms. For instance, generalization to “Lan” is notably easier on both NYUDv2 and VGGSound-S. In the Weak MG setting, existing perceptors are incorporated for their joint embedding spaces, which are typically designed for semantic purposes (e.g., CLIP (Radford et al., 2021) leverages natural language to reference learned visual concepts, enhancing the development of recent modality-binding methods (Girdhar et al., 2023; Zhu et al., 2024; Lyu et al., 2024)).
6.2. Results of Strong Modality Generalization
@@
We present the evaluation results under the strong modality generalization setting in Figures 68, and 8. Unlike the weak modality generalization setting, the testing perceptors are exclusively trained on data from novel modalities by self-supervised learning. From these results, we derive the following observations.
@@
$\bullet$ Narrowed performance gap between MML and DG. In the strong modality generalization setting, the performance gap between MML and DG methods diminishes, with MML slightly outperforming DG, especially on the NYUDv2 dataset. This contrasts with the previous weak setting, where DG methods hold a significant advantage over MML, as shown in Figure 6. Notably, MML demonstrates more robust performance across various hyperparameter searches and model selection strategies. This can be attributed to its learning objective that focuses on exploiting the complementarity among multi-modalities, not aligning well with the modality generalization goal. As a comparison, DG focuses on domain-invariant features (i.e., modal-invariant in our case), which are more sensitive when testing modalities are completely isolated.
@@
$\bullet$ Model selection remains crucial. The performance in the strong modality generalization setting still varies with different model selection methods, highlighting the importance of model selection in Figure 8. Without a validation set that mirrors the test modality, hyperparameter tuning and performance benchmarking become particularly challenging. Among the three methods, the training-modality validation set leads to more stable performance across algorithms compared to the other two model selections. The phenomenon is different from that in the case of weak modality generalization. This inspires researchers to design task-specific model selection methods or more general algorithms that can consistently perform well across multiple modality generalization approaches.
@@
$\bullet$ Limited utility of modality binding. Modality binding methods improve the generalization in weak modality generalization, while exhibiting limited utility in a more challenging task, strong modality generalization. The alignment between the training and testing perceptors helps the model preserve the relationship between modalities, leading to better consistency and accuracy when processing novel modalities in weak modality generalization testing. However, their utility diminishes in strong modality generalization, where the testing modality’s perceptor is isolated and does not share alignment with training modalities. This is due to their reliance on a joint embedding space, which struggles to generalize to unseen modalities with isolated perceptors. These findings suggest a need for alternative strategies that move beyond binding-centric paradigms to effectively handle Strong MG scenarios.
@@
$\bullet$ Narrowed generalization gap between modalities. In the Strong MG setting,the challenge of generalizing to the testing modality exhibits a more monotonic trend, as illustrated in Figure 8. The utility of modality binding, which refers to the model’s capacity to associate and leverage shared features across different modalities, is notably limited in this context. This reduction in modality binding diminishes the distinction of “easily generalizable” modalities. However, performance remains higher when “Lan” is the testing modality, particularly on NYUDv2 and VGGSound-S. Despite complete isolation of testing perceptors, connections between modalities can still be identified, which enhances generalization and substantiates our motivation to embrace Strong MG.
@@
$\bullet$ Potential of strong modality generalization. While the overall performance in the Strong MG setting is lower than that observed in the Weak MG setting, as evidenced by the comparative analysis in Figure 6), all methods achieve better-than-random predictions, indicating the ability to capture some cross-modal invariants. However, the performance gap between weak and strong modality generalization emphasizes the complexity of the strong modality generalization task, underlining the need for models that generalize without tightly bound perceptors or shared embedding spaces, such as those enhanced by self-supervised perceptors for modality generalization.
7. Discussions and Outlook
@@ 7.1. Modality Generalization Beyond Domain Generalization
Modality generalization (MG) extends beyond the traditional scope of domain generalization by focusing on enabling models to generalize across entirely unseen modalities rather than unseen domains. While domain generalization assumes similar underlying features across domains, MG introduces additional complexity due to the vast heterogeneity between modalities, such as text, audio, and visual data. Future research should emphasize developing strategies to learn and transfer invariant features across such diverse modalities. This includes moving beyond shared embedding spaces to identify deeper, task-agnostic representations that are transferable across unseen modalities. To address these challenges, researchers could explore the following aspects: (1) cross-modal invariant learning: develop mechanisms to extract invariants that are robust across modalities without assuming tight alignment or correspondence; (2) hybrid generalization frameworks: combine domain-specific techniques (e.g., meta-learning) with modality-aware methods to enhance adaptability and robustness; (3) realistic scenario validation: test models under resource-constrained or privacy-sensitive scenarios where data availability is limited for certain modalities.
7.2. Specific Algorithm Design
@@
Designing algorithms tailored to modality generalization is critical for handling the limitations of existing multi-modal learning frameworks. First, leveraging and advancing self-supervised learning to create perceptors for unseen modalities can enable effective feature extraction even without prior alignments. Second, exploring dynamic modality binding mechanisms that adapt to the availability of training modalities and the characteristics of unseen testing modalities can enhance model flexibility. Finally, developing task-aware modality generalization algorithms that integrate task-specific priors can improve generalization for downstream applications.
7.3. Construction of Related Datasets
@@
The lack of comprehensive datasets for MG research remains a major challenge. Future datasets should include diverse modalities, such as text, audio, video, depth, and bio-signals, to better simulate the complexity of real-world scenarios. These datasets should be designed to reflect MG-specific challenges, intentionally excluding certain modalities during training while ensuring their presence in the testing set to create both weak and strong MG scenarios. To facilitate benchmarking for large-scale models, datasets should also vary in complexity, modality richness, and data volume, addressing scalability concerns. Additionally, real-world constraints such as privacy and resource limitations should be considered, incorporating sensitive modalities like medical imaging or encrypted communications to emphasize the practicality of MG use cases.
7.4. Model Selection Development